OpenStudy (something):

Tiffany kicks a soccer ball off the ground and in the air with an initial velocity of 28 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 15.6 feet 14.1 feet 13.5 feet 12.3 feet

OpenStudy (lalaly):

Maximum height occurs when The derivative of H(t)=0

OpenStudy (lalaly):

so H'(t)=-32t+28 0=-32t+28 32t=28 t=0.875 H(0.875)=-16(0.875)^2+28(0.875)+0